# Uniform pathwise stability of additive singular SDEs driven by fractional Brownian motion

### v1

@@ Abstract
We study the long-time behaviour of solutions to a class of $d$-dimensional stochastic differential equations driven by fractional Brownian motion with Hurst parameter $H\in(0,1)$. The drift consists of a dissipative Lipschitz term and a singular term of regularity $\gamma>1-1/(2H)$ in Besov-Hölder scales. Well-posedness and existence of an invariant measure are established. If the singular contribution is sufficiently small, we prove that the invariant measure is unique. Our methods rely on uniform pathwise estimates which combine together the dissipativity of the drift and the regularisation effect of the noise.
@@ 1 Introduction
Our main purpose in the current work is to gain better understanding of the interlay between the different natures of the two drifts. In particular, we focus on well-posedness and the long-time behaviour of solutions to (1.1). Our main results can be summarised in a simplified form as follows.
@@
(I) If $\gamma>1/2-1/(2H)$ and $F$ has linear growth, then there exists a weak solution to the SDE (1.1). Moreover, if $\gamma>1-1/(2H)$ and $F$ is Lipschitz continuous, then there exists a unique solution to the SDE (1.1).
@@
In Section 3 and Section 4.2, we give detailed versions of the above Theorem and their proofs.
@@
In the absence of one of the drifts, equation (1.1) has been studied extensively in the literature. When $F=0$, due to the regularisation effect from the noise, equation (1.1) is well-posed even for non-Lipschitz drift $b$ (i.e. $\gamma<1$). In fact, first results for Brownian SDEs can be traced back to Zvonkin [42] and Veretennikov [40] who considered well-posedness of such SDEs with bounded measurable drifts. Their results were subsequently extended by Krylov and Röckner [24] for SDEs with integrable drifts under the so-called subcritical Ladyzhenskaya-Prodi-Serrin condition. Weak well-posedness for Brownian SDEs with distributional drifts have been considered in [33, 16, 2, 10, 14, 27, 26]. SDEs with irregular drifts driven by fractional Brownian motion also received considerable attention, with early contributions by Nualart and Ouknine [31]. [35]. The emergence of new methods such as sewing techniques have ignited renewed interests on such equations, resulting in a host of new results. Starting from Catellier and Gubinelli in [35], existence and uniqueness of strong solutions for a new class of equations have been established. We refer to [17] and the references therein for the most recent advancements. The emerging techniques also allow to identify the probability distribution of the solutions (i.e. weak well-posedness), see [6].
@@
When $b=0$ and $B$ is a Brownian motion, well-posedness and long time behaviour of equation (1.1) are well-understood, see [30] for an exhaustive analysis. For such equations, it is typically expected that the invariant measure exists uniquely and that the convergence of the law of solutions towards the stationary state is exponential (in the total variation norm). When $B$ is a fractional Brownian motion, due to long range correlations, solutions to (1.1) do not generate a Markov process. To study the long time behaviour of this class of non-Markovian processes, a prominent general framework called stochastic dynamical system has been introduced by Hairer [19]. Therein, the author established the existence and uniqueness of invariant measures, and the convergence rate of the law of an arbitrary solution towards the stationary state. This framework has been developed further in [22, 20, 23] and used for example for proving estimates on the density of the stationary state (see [28] and the references therein) and for approximating the invariant measure (see [8, 34] and references therein).
@@
Few works consider a general framework that combines both a dissipative drift and a singular drift, with the exception of [4] and [21]. In the Ph.D. thesis of the second author (see [21, Chapter 6]), a uniform-in-time bound on the moments of the solution was established for the stochastic heat equation with a distributional drift and a Lipschitz dissipative drift. This bound was for $L_{q}$-norms in space of the solutions which was not enough to deduce existence of an invariant measure in that case. In [4], the authors consider the case when $B$ is a Brownian motion, $F$ is locally Lipschitz and weakly dissipative, and $b$ is Hölder continuous. They establish existence and uniqueness of an invariant measure and an exponential contraction in the $1$-Wasserstein distance which helps them establish limit theorems. Compared to [4], we work in a general setting where $B$ is a fractional Brownian motion and $b$ can be a distribution (for small $H$). We require stronger assumptions in exchange for a stronger result (decay in moments).
@@
$|X_{t}-U_{t}|^{2}$ $\leq C\Big(1+\int_{0}^{t}e^{-\kappa_{1}(t-r)}|U_{r}|^{2}dr+\int_{0}^{t}e^{-\kappa_{1}(t-r)}\langle X_{r}-U_{r},b\left(X_{r}\right)\rangle dr\Big).$
@@
$|X_{t}-Y_{t}|^{2}\leq e^{-2\kappa_{1}t}|X_{0}-Y_{0}|^{2}+\int_{0}^{t}e^{-2\kappa_{1}(t-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr.$
@@
In Section 4, we study the existence and uniqueness of invariant measures associated to (1.1). We consider an SDE where the history of the fBm is fixed to be equal to a deterministic path and use it to define an invariant measure. We state the main results which include well-posedness of the aforementioned SDE, a link with the original SDE (the intuition being that by plugging the history of the fBm, one recovers (1.1)) existence and uniqueness of the invariant measure. The last two results are proved respectively in Section 4.1 and Section 4.2.
@@ 2 Notation, definitions and assumptions
On a probability space $(\Omega,\mathcal{F},\mathbb{P})$, we denote by $\mathbb{F}=(\mathcal{F}_{t})_{t\in\mathbb{R}}$ a filtration that satisfies the usual conditions. The conditional expectation given $\mathcal{F}_{t}$ is denoted by $\mathbb{E}^{t}$ when there is no risk of confusion on the underlying filtration. The law of a random variable $X$ is denoted by $\mathcal{L}(X)$. For any $m\in[1,\infty]$, the $L_{m}(\Omega)$ norm of $X$ is denoted by $\|X\|_{L_{m}}$ and the space $L_{m}(\Omega)$ is simply denoted by $L_{m}$.
@@
For an interval $I$ of $\mathbb{R}$, we define the simplex $\Delta_{I}=\{(s,t)\in I^{2}:s<t\}$. Moreover, for any $m\in[2,\infty)$ and $q\in[2,\infty]$, we introduce the following notation for any random variable $X$
@@
A typical example of such sequence is $(b^{k})=(P_{1/k}b)$. In the paper, we work under the Catellier-Gubinelli condition $\gamma>1-1/(2H)$, so as in the definition above, we will assume without any loss of generality that $b\in\mathcal{C}^{\gamma+}$. In fact, if $b$ is in $\mathcal{C}^{\gamma}$ then, for any $\varepsilon>0$, $b$ is in $\mathcal{C}^{(\gamma-\varepsilon)+}$.
@@
We make the following assumption on $b$:
Assumption 2.4.
$\|b\|_{\mathcal{C}^{\gamma}}\leq\Xi.$
@@ 3 Well-posedness and regularity of the solution
Let $\gamma\in(1/2-1/(2H),1)$. Assume that $F,b$ satisfy Assumptions 2.2, 2.4, respectively. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$ and $t_{0}<T$. Then there exists a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and an $\mathbb{F}$-Wiener process $W$ such that there exists a solution $(X_{t})_{t\in[t_{0},T]}$ to (1.1) with respect to $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and $W$.
@@
Moreover, $X$ satisfies (3.2) and (3.3). Furthermore, assume that $F$ satisfies Assumption 2.2 $(i)$, then $X$ also satisfies (3.1), and if $\kappa_{2}=0$, then there exists a constant $\mathbf{M}:=\mathbf{M}(m,\gamma,H,L_{F},\Xi)$ such that for
@@
$\|X_{t}-Y_{t}\|_{L_{m}}\leq|x-y|e^{\beta(t-t_{0})},\,\forall t\geq t_{0}.$ (3.6)
In the particular case where $\mathbb{F}$ is the filtration generated by $W$, one obtains a unique solution that is adapted to $\mathbb{F}$. The class of uniqueness defined in Theorem 3.2 is the same class where uniqueness is stated when $F=0$, see for example [18]. The proof of Theorem 3.2 is established in Section 3.4. The constant $\mathbf{M}$ can be tracked to Proposition 3.12 and can be explicitly computed.
@@ 3.1 Regularisation properties
For $\gamma<0$, in [18, Proposition 4.5(a)(a)], the same estimate is given for $\int f(\psi_{r}+B_{r})dr$ instead of $\int f(\psi_{r}+\widetilde{B}_{r})dr$. Hence, one can apply their result for the process $\psi-\bar{B}^{t_{0}}$, noting that $(\psi-\bar{B}^{t_{0}})+B=\psi-\widetilde{B}$ and $\llbracket\psi-\bar{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}=\llbracket\psi\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}$ to obtain (3.7). For $\gamma\geq 0$, the result follows trivially by using the supremum norm of $f$. ∎
@@
In [18, Proposition 4.7], a similar estimate is given for $\int f(\psi_{r}+B_{r})-f(\phi_{r}+B_{r})dr$. Here, the proof follows the same lines but applied for the processes $\psi-\bar{B}^{t_{0}}$ and $\phi-\bar{B}^{t_{0}}$. The differences are: we keep track of the two powers in (ts)(t-s) when applying the stochastic sewing lemma and we state the result with the semi-norm (2.6) instead of the classical Hölder norm. Noting that $(\cdot-\bar{B}^{t_{0}})+B=\cdot-\widetilde{B}$ and $\llbracket\cdot-\bar{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}=\llbracket\cdot\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}$, one can check that the proof goes through. ∎
@@ 3.2 A priori estimates
We assume that $b\in\mathcal{C}^{\infty}$ and prove estimates on the solution to (1.1) that depend only on the Besov-Hölder norm of $b$.
@@
Let $\gamma\in(1/2-1/2H,1)$. Let $F,b$ satisfy Assumptions 2.2 and 2.4, and further assume that $b\in\mathcal{C}^{\infty}$. Let $x\in\mathbb{R}^{d}$, $t_{0}\in\mathbb{R}$ and let $(X_{t})_{t\in[t_{0},\infty)}$ be the unique solution to (1.1). Let $m\in[2,\infty)$. There exists a constant $C:=C(m,\gamma,H,\kappa,\Xi)$ such that for all $t_{0}\leq s<t$ with $t-s\leq 1$, we have
@@
Let $\gamma\in(1/2-1/(2H),1)$. Let $F,b$ satisfying Assumptions 2.2 and 2.4, and further assume that $b\in\mathcal{C}^{\infty}$. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$ and $(X_{t})_{t\in[t_{0},T]}$ be the unique solution to (1.1). For any $m\in[2,\infty)$, there exists a constant $C:=C(m,\gamma,H,x,\kappa,\Xi)$ such that
@@
Let $\gamma\in(1/2-1/(2H),1)$. Let $b$ satisfy Assumption 2.4, and further assume that $b\in\mathcal{C}^{\infty}$. Let $F$ satisfy Assumption 2.3. Let $x\in\mathbb{R}^{d}$, $t_{0}\in\mathbb{R}$ and $(X_{t})_{t\in[t_{0},\infty)}$ be the solution to (1.1). Let $m\in[2,\infty)$. There exists a constant $C:=C(m,\gamma,H,L_{F},\Xi)$ such that for all $t_{0}\leq S<T$ with $T-S\leq 1$, we have
@@
Let $\gamma\in(1-1/(2H),1)$. Let $b,g$ satisfy Assumption 2.4 with the same constant $\Xi$ and let $F$ satisfy Assumption 2.3. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$ and $t_{0}<T$. Let $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ be two solutions to the SDE with the same initial condition $x$ and respective drifts $b,g$. Let $m\in[2,\infty)$ and assume that $X$ and $Y$ belong to the class $\mathcal{V}(m)$. Then there exist positive constants $C_{1}:=C_{1}(m,\gamma,H,L_{F},\Xi,T-t_{0})$ and $C_{2}=C_{2}(m,\gamma,H,\Xi)$ such that
@@
Let $\gamma\in(1-1/(2H),1)$. Let $b$ satisfy Assumption 2.4 and further assume that $b\in\mathcal{C}^{\infty}$. Let $F$ satisfy Assumption 2.3. Let $x,y\in\mathbb{R}^{d}$, $T\in\mathbb{R}$ and $t_{0}<T$. Let $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ denote two solutions to the SDE (1.1) with the same drift $b$ and respective initial conditions $x$ and $y$. Let $m\in[2,\infty)$ and assume that $X$ and $Y$ belong to the class $\mathcal{V}(m)$. There exists a constant $C:=C(m,\gamma,H,L_{F},\Xi)$ such that for any $t_{0}\leq s<t$ with $t-s\leq 1$
@@
Let $\gamma\in(1-1/(2H),1)$. Let $b$ satisfy Assumption 2.4. Let $F$ satisfy Assumption 2.3, and further assume that $F$ satisfies (2.11) with $\kappa_{2}=0$. Let $x,y\in\mathbb{R}^{d}$, $T\in\mathbb{R}$ and $t_{0}<T$. Let $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ denote two solutions to the SDE with the same drift $b$ and respective initial conditions $x$ and $y$. Let $m\in[2,\infty)$ and assume that $X$ and $Y$ belong to the class $\mathcal{V}(m)$, and assume that $X$ satisfies: there exists a constant $C_{0}$ such that for any $t_{0}<s<t$ with $t-s\leq 1$
$\llbracket X-B\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m,\infty}}\leq C_{0}.$ (3.34)
Then, there exists a positive constant $\mathbf{M}:=\mathbf{M}(m,\gamma,H,L_{F},\Xi,C_{0})$ such that for all $t\in[t_{0},T]$, we have
$\begin{split}\|X_{t}-Y_{t}\|_{L_{2m}}&\leq|x-y|\exp\{-\kappa_{1}(t-t_{0})+\mathbf{M}\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\}.\end{split}$ (3.35)

### v2

@@ Abstract
We study the long-time behaviour of solutions to a class of $d$-dimensional stochastic differential equations driven by fractional Brownian motion with Hurst parameter $H\in(0,1)$. The drift consists of a dissipative Lipschitz term and a singular term of regularity $\gamma>1-1/(2H)$ in Besov-Hölder scales. We establish well-posedness and, through a Markovian enhancement, existence of an invariant measure. If the singular contribution is sufficiently small, we prove exponential contraction of solutions, and thereby, uniqueness of the invariant measure. Our methods rely on uniform pathwise estimates which utilise together the dissipativity of the drift and the regularisation effect of the noise.
@@ 1 Introduction
Our main purpose in the current work is to gain better understanding of the interlay between the different natures of the two drifts. In particular, we focus on well-posedness and the long-time behaviour of solutions to (1.1). Our main results, which are detailed in Sections 3 and 4, can be summarised in a simplified form as follows.
@@
(I) If $\gamma>1/2-1/(2H)$ and $F$ has linear growth, then there exists a weak solution to the SDE (1.1). Moreover, if $\gamma>1-1/(2H)$ and $F$ is Lipschitz continuous, then there exists a unique strong solution to the SDE (1.1).
@@
In the absence of one of the drifts, equation (1.1) has been studied extensively in the literature. When $F=0$, due to the regularisation effect from the noise, equation (1.1) is well-posed even for non-Lipschitz drift $b$ (i.e. $\gamma<1$). In fact, first results for Brownian SDEs can be traced back to Zvonkin [43] and Veretennikov [41] who considered well-posedness of such SDEs with bounded measurable drifts. Their results were subsequently extended by Krylov and Röckner [26] for SDEs with integrable drifts under the so-called subcritical Ladyzhenskaya-Prodi-Serrin condition. Weak well-posedness for Brownian SDEs with distributional drifts have been considered in [35, 18, 2, 11, 16, 29, 28]. SDEs with irregular drifts driven by fractional Brownian motion also received considerable attention, with early contributions by Nualart and Ouknine [33]. The emergence of new methods such as sewing techniques have ignited renewed interests on such equations, resulting in a host of new results. Starting from Catellier and Gubinelli in [8], existence and uniqueness of strong solutions for a new class of equations have been established. We refer to [19] and the references therein for the most recent advancements. The emerging techniques also allow to identify the probability distribution of the solutions (i.e. weak well-posedness), see [6].
@@
When $b=0$ and $B$ is a Brownian motion, well-posedness and long time behaviour of equation (1.1) are well-understood, see [32] for an exhaustive analysis. For such equations, it is typically expected that the invariant measure exists uniquely and that the convergence of the law of solutions towards the stationary state is exponential (in the total variation norm). When $B$ is a fractional Brownian motion, due to long range correlations, solutions to (1.1) are not Markov processes. To study the long time behaviour of this class of non-Markovian processes, a prominent general framework called stochastic dynamical system has been introduced by Hairer [21]. Therein, the author established the existence and uniqueness of invariant measures, and the convergence rate of the law of an arbitrary solution towards the stationary state. This framework has been developed further in [24, 22, 25] and used for proving estimates on the density of the stationary state (see [30] and the references therein) and for approximating the invariant measure (see [9, 36] and references therein).
@@
Few works consider a general framework that combines both a dissipative drift and a singular drift, with the exception of [4] and [23]. In the Ph.D. thesis of the second author (see [23, Chapter 6]), a uniform-in-time bound on the moments of the solution was established for the stochastic heat equation with a distributional drift and a Lipschitz dissipative drift. This bound was for $L_{q}$-norms in space of the solutions which was not enough to deduce existence of an invariant measure in that case. In [4], the authors consider the case when $B$ is a Brownian motion, $F$ is locally Lipschitz and weakly dissipative, and $b$ is Hölder continuous. They establish existence and uniqueness of an invariant measure and an exponential contraction in the $1$-Wasserstein distance which helps them establish limit theorems. Compared to [4], we work in a general setting where $B$ is a fractional Brownian motion and $b$ can be a distribution (for small $H$). We require stronger assumptions in exchange for a stronger result (decay in moments).
@@
$|X_{t}-Y_{t}|^{2}\leq e^{-2\kappa_{1}(t-t_{0})}|X_{t_{0}}-Y_{t_{0}}|^{2}+\int_{t_{0}}^{t}e^{-2\kappa_{1}(t-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr.$
@@
Applications.
Our results and proofs can be useful in two main directions.
Long-time numerical approximation. Numerical schemes for (1.1) with $F=0$ have been studied for example by [23] in finite time, but understanding their long-time behaviour in the presence of $F$ and for $\gamma<0$ remains largely open. Our analysis of the invariant measure and exponential moment decay provides a natural theoretical foundation for studying the ergodic properties of discrete-time approximations. In particular, if $(b^{n})_{n\in\mathbb{N}}$ is a smooth sequence that converges to $b$ in the Besov-Hölder space of regularity $\gamma$, one might consider (as in [23]) the following tamed scheme for $h\in(0,1)$,
$X^{n,h}_{t}=x+\int_{t_{0}}^{t}F(X^{n,h}_{r_{h}})dr+\int_{t_{0}}^{t}b(X^{n,h}_{r_{h}})dr+B_{t}-B_{t_{0}},\,t\geq t_{0},$ (1.5)
where $r_{h}=h\lfloor r/h\rfloor$. Analysing the long-time properties of $X^{n,h}$ would then require proving new regularisation properties of the discretised fBm that are similar to the ones presented in this paper for the fBm. The goal here is to obtain a uniform-in-time bound on the error between $X^{n,h}$ and $X$. This would also lead to an approximation of the invariant measure, and extend to the singular regime the long-time approximation results known for smooth $b$.
Parameter estimation from invariant measures. A second natural application concerns statistical inference. Since the invariant measure characterises the long-time law of the process, it can be used to estimate a model parameter $\theta_{0}$, for instance a coefficient appearing in the drift term $(F+b)$. Given long-time discrete observations $(X_{t_{k}})_{k=0,\cdots,n}$, one can define an empirical random measure $\mu_{N}:=\frac{1}{N}\sum_{0}^{N}\delta_{X_{t_{k}}}$ and infer the unknown parameter $\theta_{0}$ by minimising a suitable distance between $\mu_{N}$ and the theoretical invariant measure $\mu_{\theta}$ (where $\theta$ denotes a general parameter). Moreover, in cases where the invariant measure is unknown, $\mu_{\theta}$ can be approximated using the scheme (1.5) enabling a fully implementable estimation procedure. We refer to [36] and references therein for the case $b=0$ but the general case still needs to be developed and would require ergodic results which our paper provides a rigorous basis for.
@@
In Section 4, we study the existence and uniqueness of invariant measures associated to (1.1). We consider an SDE where the history of the fBm is fixed to be equal to an $\mathcal{F}_{t_{0}}$ deterministic path and use it to define a Markov evolution which consists of coupling the solution with the evolution of the history of the fBm. We state the main results which include well-posedness of the aforementioned SDE, a proof of the Markov property, and existence and uniqueness of the invariant measure. The last two results are proved respectively in Section 4.1 and Section 4.2.
@@ 2 Notation, definitions and assumptions
Let $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{t\in\mathbb{R}},\mathbb{P})$ be a filtered probability space which satisfies the usual conditions. The conditional expectation given $\mathcal{F}_{t}$ is denoted by $\mathbb{E}^{t}$ when there is no risk of confusion on the underlying filtration. The law of a random variable $X$ is denoted by $\mathcal{L}(X)$. For any $m\in[1,\infty]$, the $L_{m}(\Omega)$ norm of $X$ is denoted by $\|X\|_{L_{m}}$ and the space $L_{m}(\Omega)$ is simply denoted by $L_{m}$.
@@
For an interval $I$ of $\mathbb{R}$, we denote by $\mathcal{C}(I,\mathbb{R}^{d})$ the space of continuous functions $f:I\rightarrow\mathbb{R}^{d}$ equipped with the metric
$d(f,g)=\sum_{k}2^{-k}\Big(1\wedge\sup_{r\in[-2^{k},2^{k}]\cap I}|f(r)-g(r)|\Big),$ (2.4)
and define the simplex $\Delta_{I}=\{(s,t)\in I^{2}:s<t\}$. Moreover, for any $m\in[2,\infty)$ and $q\in[2,\infty]$, we introduce the following notation for any random variable $X$
@@
A typical example of such sequence is $(b^{k})=(P_{1/k}b)$. In the paper, we work under the Catellier-Gubinelli condition $\gamma>1-1/(2H)$, so as in the definition above, we will assume without any loss of generality that $b\in\mathcal{C}^{\gamma+}$. In fact, if $b$ is in $\mathcal{C}^{\gamma}$ then, for any $\varepsilon>0$, $b$ is in $\mathcal{C}^{(\gamma-\varepsilon)+}$. Therefore, we make the following assumption on $b$:
Assumption 2.2.
$\|b\|_{\mathcal{C}^{\gamma}}\leq\Xi.$
@@ 3 Well-posedness and regularity of the solution
Let $\gamma\in(1/2-1/(2H),1)$. Assume that $F,b$ satisfy Assumptions 2.3, 2.2, respectively. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$ and $t_{0}<T$. Then there exists a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and an $\mathbb{F}$-Wiener process $W$ such that there exists a solution $(X_{t})_{t\in[t_{0},T]}$ to (1.1) with respect to $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and $W$.
@@
Moreover, $X$ satisfies (3.2) and (3.3). Furthermore, let assume that $F$ satisfies Assumption 2.3, then $X$ also satisfies (3.1), and if $\kappa_{2}=0$, then for any $m\in[2,\infty)$, there exists a universal constant $C$ and a constant $\mathbf{M}:=\mathbf{M}(m,\gamma,H,L_{F},\Xi)$ such that for
@@
$\|X_{t}-Y_{t}\|_{L_{m}}\leq C|x-y|e^{\beta(t-t_{0})},\,\forall t\geq t_{0}.$ (3.6)
In the particular case where $\mathbb{F}$ is the filtration generated by $W$, one obtains a unique solution that is adapted to $\mathbb{F}$. The class of uniqueness defined in Theorem 3.2 is the same class where uniqueness is stated when $F=0$, see for example [20]. The proof of Theorem 3.2 is established in Section 3.4. The universal constant $C$ and the constant $\mathbf{M}$ can be tracked to the proof of Proposition 3.12 and can be explicitly computed.
@@ 3.1 Regularisation properties
For $\gamma<0$, in [20, Proposition 4.5(a)(a)], the same estimate is given for $\int f(\psi_{r}+B_{r})dr$ instead of $\int f(\psi_{r}+\widetilde{B}_{r})dr$. Hence, one can apply their result for the process $\psi-\bar{B}^{t_{0}}$, noting that $(\psi-\bar{B}^{t_{0}})+B=\psi-\widetilde{B}$ and $\llbracket\psi-\bar{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}=\llbracket\psi\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}$ to obtain (3.7). For $\gamma\geq 0$, the result follows trivially by using the supremum norm of $f$. ∎
@@
In [20, Proposition 4.7], a similar estimate is given for $\int f(\psi_{r}+B_{r})-f(\phi_{r}+B_{r})dr$. Here, the proof follows the same lines but applied for the processes $\psi-\bar{B}^{t_{0}}$ and $\phi-\bar{B}^{t_{0}}$. The differences are: we keep track of the two powers in (ts)(t-s) when applying the stochastic sewing lemma and we state the result with the semi-norm (2.7) instead of the classical Hölder norm. Noting that $(\cdot-\bar{B}^{t_{0}})+B=\cdot-\widetilde{B}$ and $\llbracket\cdot-\bar{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}=\llbracket\cdot\rrbracket_{\mathcal{C}^{\tau}_{[s,t]}L_{m,q}}$, one can check that the proof goes through. ∎
@@ 3.2 A priori estimates
We assume that $b\in\mathcal{C}^{\infty}$. For any $x\in\mathbb{R}^{d}$ and $t_{0}\in\mathbb{R}$, we write $(X_{t})_{t\geq t_{0}}$ for the solution of (1.1) starting from $x$ at time $t_{0}$, and prove estimates on $X$ that depend only on the Besov-Hölder norm of $b$.
@@
Let $\gamma\in(1/2-1/2H,1)$ and $m\in[2,\infty)$. Let $F,b$ satisfy Assumptions 2.3 and 2.2, and further assume that $b\in\mathcal{C}^{\infty}$. There exists a constant $C:=C(m,\gamma,H,\kappa,\Xi)$ such that for any $x\in\mathbb{R}^{d}$, any $t_{0}\in\mathbb{R}$, and any $t_{0}\leq s<t$ with $t-s\leq 1$, we have
@@
Let $\gamma\in(1/2-1/(2H),1)$ and $m\in[2,\infty)$. Let $F,b$ satisfying Assumptions 2.3 and 2.2, and further assume that $b\in\mathcal{C}^{\infty}$. There exists a constant $C:=C(m,\gamma,H,x,\kappa,\Xi)$, such that for any $x\in\mathbb{R}^{d}$, any $T\in\mathbb{R}$, any $t_{0}<T$, we have
@@
Let $\gamma\in(1/2-1/(2H),1)$. Let $b$ satisfy Assumption 2.2, and further assume that $b\in\mathcal{C}^{\infty}$. Let $F$ satisfy Assumption 2.4. There exists a constant $C:=C(m,\gamma,H,L_{F},\Xi)$ such that for any $x\in\mathbb{R}^{d}$, any $t_{0}\in\mathbb{R}$, any $t_{0}\leq S<T$ with $T-S\leq 1$, we have
@@
Let $\gamma\in(1-1/(2H),1)$, $m\in[2,\infty)$, $T\in\mathbb{R}$ and $t_{0}<T$. Let $b,g$ satisfy Assumption 2.2 with the same constant $\Xi$ and let $F$ satisfy Assumption 2.4. There exist positive constants $C_{1}:=C_{1}(m,\gamma,H,L_{F},\Xi,T-t_{0})$ and $C_{2}=C_{2}(m,\gamma,H,\Xi)$ such that for any $x\in\mathbb{R}^{d}$, any two solutions $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ to the SDE (1.1) with the same initial condition $x$ and respective drifts $b,g$ such that $X$ and $Y$ belong to the class $\mathcal{V}(m)$, we have
@@
Let $\gamma\in(1-1/(2H),1)$ and $m\in[2,\infty)$. Let $b$ satisfy Assumption 2.2 and further assume that $b\in\mathcal{C}^{\infty}$. Let $F$ satisfy Assumption 2.4. There exists a constant $C:=C(m,\gamma,H,L_{F},\Xi)$ such that for any $x,y\in\mathbb{R}^{d}$, any $T\in\mathbb{R}$, any $t_{0}<T$, any two solutions $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ to the SDE (1.1) with the same drift $b$ and respective initial conditions $x$ and $y$ such that $X$ and $Y$ belong to the class $\mathcal{V}(m)$, and any $t_{0}\leq s<t$ with $t-s\leq 1$, we have
@@
Let $\gamma\in(1-1/(2H),1)$ and $m\in[2,\infty)$. Let $b$ satisfy Assumption 2.2. Let $F$ satisfy Assumption 2.4, and further assume that $F$ satisfies (2.12) with $\kappa_{2}=0$. There exists a universal constant $C$ and a positive constant $\mathbf{M}:=\mathbf{M}(m,\gamma,H,L_{F},\Xi,C_{0})$ such that for any $x,y\in\mathbb{R}^{d}$, any $T\in\mathbb{R}$ any $t_{0}<T$, any two solutions $(X_{t})_{t\in[t_{0},T]},(Y_{t})_{t\in[t_{0},T]}$ the SDE (1.1) with the same drift $b$ and respective initial conditions $x$ and $y$ such that $X$ and $Y$ belong to the class $\mathcal{V}(m)$ and $X$ satisfies: there exists a constant $C_{0}$ such that for any $t_{0}<s<t$ with $t-s\leq 1$
$\llbracket X-B\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m,\infty}}\leq C_{0},$ (3.34)
any $t\in[t_{0},T]$, we have
$\begin{split}\|X_{t}-Y_{t}\|_{L_{2m}}&\leq C|x-y|\exp\{-\kappa_{1}(t-t_{0})+\mathbf{M}\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\}.\end{split}$ (3.35)